OpenStudy (anonymous):

WILL GIVE MEDAL Find an equation for the nth term of the sequence. -3, -12, -48, -192, .

OpenStudy (freckles):

hint: it is a geometric sequence I know this because it has a common ratio

OpenStudy (anonymous):

what?

OpenStudy (freckles):

common ratio exist when you have term divided by previous term is equal to some number constantly

OpenStudy (freckles):

that is; \[\frac{a_2}{a_1}=\frac{a_3}{a_2}=\frac{a_4}{a_3}= \cdots =\frac{a_n}{a_{n-1}}=r\] r is called the common ratio

OpenStudy (freckles):

you should see the following for yours... \[\frac{-12}{-3}=\frac{-48}{-12}=\frac{-192}{-48}=?\]

OpenStudy (freckles):

\[\text{ geometric sequence } \\ a,ar,ar^2,ar^3, \cdots ,ar^{n-1}, \cdots \\ a_1=a \\ a_2=ar \\ a_3=ar^2 \\ \cdots \\ a_n=ar^{n-1}\] so you just need to determine r and the first term a from above

OpenStudy (freckles):

and plug into that nth term I have above